OpenStudy (princeevee):

solve 1= -(-w+6) 3 -3 7 -7

OpenStudy (anonymous):

you can solve this or you can plug in the numbers to see which one works for example 3 does not work because \[-(-3+6)=-3\neq 1\]

OpenStudy (princeevee):

so its b?

OpenStudy (anonymous):

it is usually C, try that oen

OpenStudy (princeevee):

so the answer is c?

OpenStudy (anonymous):

yes

OpenStudy (princeevee):

thanks

OpenStudy (princeevee):

can you help me with 3 more?

OpenStudy (anonymous):

ok

OpenStudy (princeevee):

solve for x lxl - 5 = 7 x = -12 or 12 x=12 x=-2 or 2 x=2

OpenStudy (anonymous):

x-5=7 or -x-5=7

OpenStudy (anonymous):

x=7+5 and x=(7+5)/-1

OpenStudy (princeevee):

but what is x?

OpenStudy (anonymous):

hmmm lets see if we can do this

OpenStudy (anonymous):

\[|x|-5=7\] right?

OpenStudy (princeevee):

yes

OpenStudy (anonymous):

add 5 to both sides, to get rid of the \(-5\) on the left what do you get?

OpenStudy (princeevee):

i still dont get but i think d?

OpenStudy (princeevee):

@satellite73